Use a table to find the break-even point. Check your solution.
C=15x + 150
R=45x

Answer by stanbon(75887)

find the break-even point. Check your solution.
C=15x + 150
R=45x
------
Solve:: C = R
15x + 150 = 45x
30x = 150
x = 5
---
The R = C = 45*5 = 225
-----------
Cheers,
Stan H.
--------------------